Help please!
For Faculty Appreciation Day, the Faculty Club Lunch Room asked each faculty member to leave a business card in a container. At the end of lunch a card would be drawn, and that faculty member would be given free lunches for a week. Eighty-eight faculty left a card, including 5 mathematics faculty, 11 history faculty, and 2 chemistry faculty. Find the probability of the following. (to three decimal places.)
(a) The card drawn was for a mathematics faculty member.
(b) The card drawn was for a chemistry faculty member.
(c) The card drawn was for a history faculty member.
(d) The card drawn was for a mathematics or chemistry faculty member.

Answer by sudhanshu_kmr(1152)


(a) The card drawn was for a mathematics faculty member.
probability = 5/88

(b) The card drawn was for a chemistry faculty member.
probability = 2/88


(c) The card drawn was for a history faculty member.
probability = 11/88 = 1/8


(d) The card drawn was for a mathematics or chemistry faculty member.
probability = 7/88